OpenStudy (racquelpellikan):

A spherical globe has a radius of 9 inches. What is the surface area of the globe to the nearest hundredth of a square inch?

OpenStudy (leader):

Welcome to OpenStudy! \(\sf\Large Surface~Area= 4\times \pi \times r^2\)

OpenStudy (racquelpellikan):

Thank you for your help.